OpenStudy (anonymous):

1. Divide (18x^3 + 12x^2 - 3x) / 6x^2 A.) x^2 + x - 1/2 x B.) 3x + 2 C.) 3x + 2 - 1/2x D.) 3x - 1/2x 2. Divide. (2x^3 - x^2 - 24x + 12) / (2x - 1) A.) x^2 - 1 B.) x^2 - 12 C.) 2x - 1 D.) x^2 - 1/2 x - 12 3. Simplify. 2a + 2/a^2 - 1 A.) 2/a - 1 B.) 2a + 2/ a^2 - 1 C.) -2 D.) a + 1/ a - 1 4. What is the undefined value of 3b^2 + 13b + 4/ b + 4? A.) 4 B.) 0 C.) -3 D.) -4 5. Multiply. x^2 - x - 42/ x - 8 (times) x^2 - 64/ 9x + 54 A.) (x + 6) (x - 7)/ 9(x + 8) B.) (x - 7) (x + 8)/ 9 C.) (x - 7) (x + 8)/ 9x + 54 D.) (x + 6) (x - 7)/ 9

OpenStudy (anonymous):

#is b

OpenStudy (anonymous):

6. Divide. x^2 + 9x + 20/x^2 - 25 (divided by) x + 4/x - 4 A.) x - 4/x - 5 B.) x + 5/ x - 5 C.) x + 5/ x - 4 D.) 9x + 4/ 5 7. Add. 8 + x - 1/ x + 4 A.) 9x + 31/ x + 4 B.) 7 + x/ x + 4 C.) 8x - 8/ 8x + 34 D.) 9x + 31/ x - 4 8. Subtract. (x^2 - 1/ x^2 - x - 2) - (x - 1/ x - 2) A.) 1 B.) 0 C.) -1 D.) x - 1/x - 2 9. Solve. 10/3x + 4/3 = 7 + x/ 2x A.) x = 1/3 B.) x = 17/5x C.) x = 1/5 D.) x = 1/6 10. x + 12/x + 4 = x/ x + 8 A.) x = -4 B.) x = 8 C.) x = 6 D.) x = -6

OpenStudy (anonymous):

? @JULIANCHICO

OpenStudy (anonymous):

go to this web site its will give u the answers Mathway.com

OpenStudy (anonymous):

11. A group of college students is volunteering for Homes for the community during their spring break. They are putting the finishing touches on a house they built. Working alone, Wade can paint a certain room in 7 hours. Rhonda can paint the same room in 5 hours. How long will it take them working together to paint the room? Round your answer to the nearest hundredth, if necessary. A.) 17.5 hours. B.) 0.08 hours. C.) 6 hours. D.) 2.92 hours.

OpenStudy (anonymous):

Can you help me with these? @JULIANCHICO

OpenStudy (anonymous):

I don't want to use that website.

OpenStudy (anonymous):

d

OpenStudy (anonymous):

hell naa bro this is too much work negga

OpenStudy (anonymous):

I didn't ask if you could do ALL of it. Just the ones you can answer.

OpenStudy (anonymous):

D for number one?

OpenStudy (anonymous):

naa my negga am too lazy anyways bro

OpenStudy (nurali):

11. SInce wade paints a room in 7 hrs, that means that w = 1/7(room/hr) SInce rhonda paints a room in 5 hrs, that means that r = 1/5(room/hr) now the amount of time that it will take them to finish a room will be 1/(the rate at which wade works + the rate at which rhonda works) this will give 2.92 hours

OpenStudy (anonymous):

Don't bro me if you don't know me. And thanks @Nurali can you help with some of the other questions? @Nurali

OpenStudy (anonymous):

naa d for #11

OpenStudy (anonymous):

post the questions on answers.yahoo.com they will defenily answer this 4 u

OpenStudy (anonymous):

Can you put them for me? Don't have yahoo.

OpenStudy (dwelsher):

c

OpenStudy (anonymous):

For number what?

OpenStudy (anonymous):

@dwelsher

OpenStudy (anonymous):

Guys I really need help. I don't want people just wandering, I need help, now. So please if you can, help. Thanks.

OpenStudy (anonymous):

What about 5, 9, and 10? @Nurali